# RadarLLM: Adapting Pretrained Large Language Models for Marine Radar Target Detection with Preference-aware Loss

### v2

@@ Abstract
Recent advances in pre-trained large language models (LLMs) have demonstrated their capacities to capture universal knowledge, making them promising general-purpose optimization solvers for wireless signal processing. Motivated by these findings, we take the first step towards fine-tuning pre-trained LLMs for the effective analysis of radar signal features in marine target detection tasks. Nevertheless, directly fine-tuning pre-trained LLMs on marine target detection tasks tends to suffer from pronounced overfitting, particularly in challenging low signal-to-clutter ratio (SCR) scenarios. This overfitting is mainly due to the model’s tendency to memorize noisy feature patterns rather than learning discriminative structures that generalize well to unseen data. To address this challenge, we introduce RadarLLM, a novel fine-tuning framework that utilizes an effective preference-aware loss. Unlike conventional training strategies that uniformly optimize all feature tokens, this loss function selectively optimizes different feature patches based on their online evaluated learning values, thus guiding the model to focus on the most generalizable patterns during optimization. We theoretically demonstrate the effectiveness of the evaluated learning values by transforming the problem as selecting useful feature tokens. Extensive experiments on real-world marine radar datasets show that 1) the proposed loss function outperforms the original one, showing particularly significant improvements under challenging low SCR conditions, with an average performance gain of 9.9% and 2) RadarLLM consistently outperforms state-of-the-art baselines in diverse detection scenarios, with particularly notable gains under limited training data conditions.
@@
Marine small target detection, radar signal processing, pre-trained large language models (LLMs), preference-aware loss.
@@ I Introduction
Despite recent progress, compact NN models with small parameter scales and randomly initialized weights often exhibit limited learning capabilities and unstable training processes. These limitations lead to poor detection performance, particularly in scenarios with scarce training data or with significant environmental variability. Extensive recent studies [18, 19, 20, 21, 22, 23, 24] have shown that pre-trained LLM can be a good starting point for wireless signal and time series modeling, as they are pre-trained on massive sequential datasets. Compared to random initialized weights of small models, this starting point may significantly improve model optimization stability and effectiveness. Moreover, the self-attention mechanism in LLM exhibits structural similarity to Principal Component Analysis (PCA) [18], demonstrating strong generality in extracting key features from high-dimensional sequence data. This insight opens a very promising direction for radar target detection by fine-tuning pre-trained LLMs, offering the potential for performance improvements.
Our preliminary study [25] shows that fine-tuning a pre-trained GPT2 model to analyze sequence features yields state-of-the-art detection performance. Extensive ablation experiments further confirm that removing the GPT2 model or substituting it with a randomly initialized Transformer encoder significantly degrades the detection rate.
@@
Although powerful and promising, some challenges still exist. First, the large scale of LLM requires lightweight and efficient fine-tuning strategies, which are essential for practical deployment, especially in resource-constrained settings. Second, and more critically, for radar applications, adapting LLMs to process sequence features in low signal-to-clutter ratio (SCR) conditions is far from a routine task, often leading to severe overfitting issues. This phenomenon is fundamentally attributed to the intrinsic characteristics of sequence features:
@@
  • Significant Clutter-Induced Noise: In low SCR environments, small targets are frequently obscured by ocean waves. This causes large noise in sequence features, especially in phase and amplitude features, increasing the likelihood of the model overfitting to irrelevant and noisy information.
@@
  • Feature Pattern Convergence Imbalance: Various patterns in different sequence features often exhibit different convergence rates during fine-tuning. For example, simple and regular patterns may rapidly reach convergence during model training, and then exhibit a tendency for overfitting, whereas other valuable patterns may require more iterations to achieve convergence.
To address the aforementioned challenges, this paper introduces RadarLLM, an innovative framework designed for learning from radar echo signals. First, we extract five sequence features from radar echo signals and patch them into multiple feature tokens. Second, we train a small reference model based on the validation samples and calculate the loss for each feature token of the training samples to estimate its generalization potential. Then, we construct a target detector based on a pre-trained LLM and fine-tune the model using a lightweight finetuning module, along with a novel preference-aware loss function inspired by the work in NLP [26] and CV [27]. This loss function introduces a selective training strategy by assigning different training weights to feature patches according to their online evaluated learning values, significantly improving model performance in low SCR environments. Finally, we retrain the classification head to enhance detection performance and control the false alarm rate.
@@
    We propose RadarLLM, a novel framework combining a lightweight reference model and a pre-trained LLM for marine radar target detection. To the best of our knowledge, we are the first to show that the well-known LLMs can be a strong starting point for intelligent radar signal processing.
@@
    For the fine-tuning of the LLMs, we propose a novel preference-aware loss function. This loss function can selectively optimize useful feature tokens with high learning values, significantly reducing the risk of model overfitting.
@@
    Our extensive experiments confirm the effectiveness of RadarLLM. It attains new state-of-the-art detection performance under sufficient training data and constrained training data settings on popular marine radar datasets.
The remainder of this paper is structured as follows. Section II briefly summarizes the related work. Section III presents the detailed operating procedure of the proposed RadarLLM. Section IV introduces the experimental results and analysis. Section V concludes our work.
@@ II-A Deep Learning Powered Marine Radar Target Detection
In addition to traditional supervised learning, semi-supervised, unsupervised, and incremental learning approaches are proposed to further advance radar target detection. Wang et al. [28] propose a self-evolving framework for maritime radar target detection using semi-supervised learning, which enhances detection performance through unsupervised sample selection, data augmentation, and model optimization. Xia et al. [10] introduce an unsupervised contrastive learning framework that learns discriminative representations between targets and clutter from unlabeled data. Wang et al. [29] propose an incremental learning-based target detection method, which can continuously adapt the NN model in real time according to environmental changes. These methods provide compelling evidence that deep learning can significantly enhance detection performance without the need for explicit mathematical modeling. Despite their initial promise, these methods are inherently constrained by their dependence on relatively compact neural networks. Such models exhibit limited representational capacity, which hinders their ability to learn complex, high-dimensional feature representations. Consequently, they are particularly susceptible to performance degradation in the face of diverse and nonstationary environmental conditions. These challenges motivate us to first explore the use of LLMs for radar target detection.
@@ II-B Language Model Powered Signal Processing Task
Recently, LLMs such as ChatGPT and DeepSeek have demonstrated exceptional capabilities in natural language understanding, code generation, math problem solving, and human-like text generation [30]. Some studies also demonstrate their potential in the signal processing (SP) task. For example, Zhou et al. [18] propose a unified time series analysis framework by fine-tuning frozen GPT2 [31] and achieving SOTA performance on various datasets. Liu et al. [24] propose a GPT2-empowered channel prediction framework to improve prediction accuracy. Sheng et al. [32] utilize the GPT2 model to develop an effective and robust beam prediction method. Zheng and Dai [21] propose a multi-task LLM framework to satisfy the requirements of different wireless communication tasks. IOT-LLM and Penetrative AI [33, 34] directly utilize raw signal data as input and chain of thought prompts for reasoning on the Internet of Things (IOT) tasks. However, recent studies [35] have raised concerns about the universal applicability of LLM in signal processing tasks. LLM-based approaches do not always deliver significant performance improvements, implying that their pre-trained parameters sometimes fail to transfer to downstream tasks. In marine radar target detection under low SCR conditions, directly fine-tuning a pre-trained LLM often leads to severe overfitting issues, limiting its parameter transfer capability. To mitigate this limitation, we introduce a preference-aware loss that selectively trains on informative feature patches.
@@ III Proposed Method
Figure 1: Overview of our RadarLLM framework.
III-A Overview of the RadarLLM Framework
In this study, we present RadarLLM, an LLM–empowered framework tailored for the marine radar target detection task. RadarLLM training protocol is illustrated in Fig. 1 and shown in Algorithm 1.
To enable lightweight yet effective fine-tuning of LLMs for radar target detection, the proposed RadarLLM framework integrates four tightly coupled modules into a unified optimization pipeline. First, the sequence feature extraction and patching module (Section III-B) derives five discriminative sequence features from the multi-domain transformation of radar echo signals. Then a patching operation is applied to capture local semantic patterns while reducing computational complexity. Second, a lightweight finetuning module (Section III-C) is introduced to efficiently adapt the LLM backbone to the radar task, preserving its universal knowledge contained while minimizing parameter updates. Third, a preference-aware loss function (Section III-D), which utilizes a lightweight reference model, enables selective training on informative feature patches and effectively alleviates overfitting problem. Finally, an autoencoder-based classification head (Section III-E) is retrained to enhance the detection rate and maintain a controllable false alarm rate.
As illustrated in Fig. 1, the LLM-based model first employs an input embedding layer to project the feature patches into the same latent space as the pre-trained LLM. The embedded representations are then processed by the initial encoder layers of the pre-trained LLM to extract more discriminative features. Subsequently, a layer normalization operation refines these features, yielding the normalized representation $F^{\mathrm{LN}}$, which enhances consistency and stability. Finally, an output linear layer generates token-level detection results. For the structure of the reference model, a standard Transformer encoder [36] is utilized to process the input feature patches and produce token-level detection outputs, as illustrated in Fig. 2.
Algorithm 1 RadarLLM Training Protocol
1:
2:Radar echo signals $\{x(n)\}_{n=1}^{N}$ and Target labels $\{y_{b}\}_{b=1}^{B}$
3:Pre-trained LLM parameters $\theta_{\text{LLM}}$
4:Stage 1: Sequence Feature Extraction
  1. 1.
  2. 2.
    Patch the five features into feature tokens, see Fig. 4.
5:Stage 2: Reference Model Training
  1. 1.
  2. 2.
6:Stage 3: Fine-tune LLM for Target Detection
  1. 1.
  2. 2.
    Fine-tune LLM for target detection via preference-aware loss function, see Eq.(14) and Eq.(16).
7:Stage 4: Classification Head Retraining
  1. 1.
    Train autoencoder classification head based on the weighted sum of reconstruction loss and CE loss, see Eq.(22), Eq.(23), and Eq.(24).
  2. 2.
    Adjust detection threshold $\eta$ for controllable FAR, see Eq.(25).
8:
9:Fine-tuned RadarLLM model with controllable FAR.
@@ III-B Sequence Feature Extraction and Patching
Figure 4: An illustration of the patching operation.
III-C Lightweight Fine-tuning Module for the LLM Backbone
Inspired by the strategy proposed in [18], we adopt a lightweight fine-tuning module that jointly updates the layer normalization parameters and integrates a Low-Rank Adaptation (LoRA) module [37]. LoRA is a parameter-efficient adaptation technique that freezes the pre-trained model weights while injecting trainable low-rank decomposition matrices into selected layers.
This design offers two key advantages. First, given the immense parameter scale of LLMs, full fine-tuning is often computationally prohibitive. This module mitigates this limitation by updating only compact low-rank matrices, thereby reducing trainable parameters by several orders of magnitude. Second, this module alleviates catastrophic forgetting of pre-trained knowledge during adaptation, as the newly acquired information is encoded within a subspace of lower rank than the original representation. Consequently, this module enables the finetuned model to retain its pre-trained general-purpose knowledge while efficiently adapting to downstream radar tasks.
Concretely, we fine-tune the query and value matrices, $\mathbf{W}_{Q}$ and $\mathbf{W}_{V}$, within each attention block of the LLM backbone, as illustrated in Fig. 5. LoRA introduces two trainable rank decomposition matrices, $A$ and $B$, attached to the frozen pre-trained weight matrix $\mathbf{W}$. The modified transformation is expressed as:
@@ III-C Lightweight Fine-tuning Module for the LLM Backbone
where $\mathbf{W}\in\mathbb{R}^{d{1}\times d{2}}$, $A\in\mathbb{R}^{d_{1}\times r}$, $B\in\mathbb{R}^{d_{2}\times r}$, and $r\ll\min(d_{1},d_{2})$. The matrices are initialized as $A\sim\mathcal{N}(0,\sigma^{2})$ and $B=0$. During fine-tuning, $\mathbf{W}$ remains frozen while $A$ and $B$ are optimized. Owing to the low-rank constraint ($r\ll\min(d_{1},d_{2})$), this approach dramatically reduces the number of trainable parameters compared to full-parameter fine-tuning.
@@ III-D1 Methodology
To enhance the fine-tuning of LLMs for radar target detection under low SCR conditions, we introduce a token-level preference-aware loss function. In such scenarios, a substantial proportion of feature patches contribute little information for effective learning. Uniform optimization over all feature patches risks overfitting by coercing the model to memorize uninformative or even misleading patterns. To mitigate this, we propose a dynamic re-weighting scheme that adaptively assigns higher learning weights to feature patches exhibiting superior generalization potential. Specifically, our approach uses a dual-model framework consisting of a lightweight reference model $\theta_{r}$ and a high-capacity target model $\theta_{t}$, which leverages LLMs.
  • Reference Model: The reference model is built with a lightweight Transformer encoder and trained solely on validation samples using standard cross-entropy loss. By leveraging the reference model, we can estimate the generalization potential of each feature token in the training samples. A lower reference loss indicates that the token can be effectively captured by a lightweight model, suggesting strong generalization potential. In contrast, a higher reference loss typically reflects excessive noise, making the token less useful for improving detection performance.
  • Target Model: A significantly larger model built upon a pre-trained LLM, parameterized by $\theta_{t}$.
For a batch of training samples, we compute the token-level cross-entropy losses for both models. Given a feature token $F_{b,k}^{\mathrm{fi}}$, the corresponding two losses are as follows:
  • Reference Loss:
@@
    where $\hat{y}_{b,k}^{i,\theta_{r}}$ is the prediction from $\theta_{r}$.
  • Target Loss:
@@
    where $\hat{y}_{b,k}^{i,\theta_{t}}$ is generated by the fine-tuned much larger model $\theta_{t}$.
The core of our method lies in computing a token importance score $s_{b,k}$ based on the discrepancy between the target loss and reference loss:
@@
where $\alpha$ is a scaling coefficient, and $\mathrm{ReLU}(\cdot)$ ensures that only tokens with positive excess loss are trained. Tokens with low importance score are not prioritized for optimization because they have already been well learned, as indicated by a low target loss $\mathcal{L}_{\theta_{{t}},b,k}$, or with large noise, as reflected by a high reference loss $\mathcal{L}_{\theta_{{r}},b,k}$. The $\mathrm{ReLU}(\cdot)$ function is defined as:
$\operatorname{ReLU}(x)=\left\{\begin{array}[]{ll}x&x>0\\ 0&x\leq 0\end{array}\right$. (15)
More detailedly, this strategy results in the following selective training scheme:
  • If the target loss is much larger than the reference loss, the token is deemed with large learning value, and is therefore prioritized for training.
  • If the reference loss is large, the token shows weak generalization potential and a high risk of overfitting; therefore, its training weight is reduced during optimization.
  • If the target loss is already small, the token has been sufficiently learned, and further training brings little gain, the training weight is also reduced to avoid redundant updates.

### v3

# RadarPLM: Adapting Pretrained Language Models for Marine Radar Target Detection with Preference-aware Loss

@@ Abstract
Recent advances in pre-trained language models (PLMs) have demonstrated their capabilities in capturing universal knowledge, making them promising applications for radar signal processing. Nevertheless, directly fine-tuning PLMs on radar signals is both computationally expensive and prone to overfitting, particularly in low signal-to-clutter ratio (SCR) environments. In this paper, we propose a novel fine-tuning framework for PLM-based marine radar target detection. First, we design a lightweight adaptation module, enabling parameter-efficient fine-tuning while preserving the pretrained model’s general knowledge. Second, a novel preference-aware loss is developed to selectively optimize different feature patches based on their online evaluated learning values, guiding the model to concentrate on the most generalizable feature patterns during optimization. Extensive experiments on real-world marine radar datasets demonstrate that the proposed finetuning framework achieves an average performance improvement of 9.9% over the standard approach under low SCR conditions. Furthermore, the fine-tuned model, RadarPLM, consistently outperforms state-of-the-art detectors, particularly when training data are limited.
@@
Marine small target detection, radar signal processing, pre-trained language models (PLMs), preference-aware loss.
@@ I Introduction
Despite recent progress, compact NN models with small parameter scales and randomly initialized weights often exhibit limited learning capabilities and unstable training processes. These limitations lead to poor detection performance, particularly in scenarios with scarce training data or with significant environmental variability. Extensive recent studies [18, 19, 20, 21, 22, 23, 24] have shown that pre-trained language models (PLMs) can be good starting point for wireless signal and time series modeling, as they are pre-trained on massive sequential datasets. Compared to random initialized weights of small models, this starting point may significantly improve model optimization stability and effectiveness. Moreover, the self-attention mechanism in PLM exhibits structural similarity to Principal Component Analysis (PCA) [18], demonstrating strong generality in extracting key features from high-dimensional sequence data. This insight opens a very promising direction for radar target detection by fine-tuning PLMs, offering the potential for performance improvements.
@@
Although powerful and promising, some challenges still exist. First, the large scale of PLM requires lightweight and efficient fine-tuning strategies, which are essential for practical deployment, especially in resource-constrained settings. Second, for practical radar applications, adapting PLMs to process sequence features in low signal-to-clutter ratio (SCR) conditions often leads to severe overfitting issues. This phenomenon is fundamentally attributed to the intrinsic characteristics of sequence features:
@@
  • Significant Clutter-Induced Noise: In low SCR environments, small targets are frequently obscured by ocean waves, which inevitably generate mixed-statue segments. The mixed statue segments can cause a low SCR and increase the likelihood of the model overfitting to irrelevant and noisy information.
@@
  • Feature Pattern Convergence Imbalance: Various patterns in different sequence features often exhibit different convergence rates during fine-tuning. For example, simple and regular patterns may rapidly reach convergence during model fine-tuning and then exhibit a tendency for overfitting, whereas other valuable patterns may require more iterations to achieve convergence.
Consequently, uniform optimization across all feature patterns may cause the model to overfit not only to noisy patterns but also to overly simplistic ones, which fail to capture the more generalizable patterns inherent in radar sequence features.
To overcome the above-mentioned issues, including the demand for lightweight fine-tuning and the tendency of overfitting under low SCR conditions, we develop a novel framework named RadarPLM. The proposed framework effectively bridges the gap between PLMs and radar target detection tasks through lightweight adaptation module and selective training. First, we extract five sequence features from radar echo signals and patch them into multiple feature tokens. Next, a radar target detector is constructed based on a PLM, which is fine-tuned through a lightweight adaptation module to efficiently align with the radar detection task. For model optimization, we design a preference-aware loss function by incorporating a lightweight reference model, which is inspired by recent advances [25, 26]. This loss introduces a selective training strategy that dynamically adjusts the learning weights of different feature patches according to their online-evaluated learning values, significantly mitigating model overfitting. Finally, the binary classification head is retrained to further improve detection performance.
Our preliminary study [27] demonstrated that fine-tuning a pre-trained GPT2 model effectively extracts discriminative information from sequence features, achieving state-of-the-art detection performance. However, model overfitting remains a challenge. To address this, we propose the RadarLLM framework, which introduces a lightweight adaptation module and selective training strategy for more reliable and generalizable adaptation.
@@
    We propose RadarPLM, a novel framework combining a lightweight reference model and a fine-tuned PLM for marine radar target detection. To the best of our knowledge, we are the first to show that the well-known LLMs can be a strong starting point for intelligent radar signal processing.
@@
    For effective fine-tuning, we develop a lightweight adaptation module together with a novel preference-aware loss function, which jointly reduce computational overhead and effectively mitigate the risk of model overfitting.
@@
    Our extensive experiments validate that RadarLLM achieves new state-of-the-art detection performance on popular marine radar datasets under both sufficient and limited training data settings.
The remainder of this paper is structured as follows. Section II briefly summarizes the related work. Section III presents the detailed operating procedure of the proposed RadarPLM. Section IV introduces the experimental results and analysis. Section V concludes our work.
@@ II-A Deep Learning Powered Marine Radar Target Detection
In addition to traditional supervised learning, semi-supervised, unsupervised, and incremental learning approaches are proposed to further advance radar target detection. Wang et al. [28] propose a self-evolving framework for maritime radar target detection using semi-supervised learning, which enhances detection performance through unsupervised sample selection, data augmentation, and model optimization. Xia et al. [10] introduce an unsupervised contrastive learning framework that learns discriminative representations between targets and clutter from unlabeled data. Wang et al. [29] propose an incremental learning-based target detection method, which can continuously adapt the NN model in real time according to environmental changes. These methods provide compelling evidence that deep learning can significantly enhance detection performance without the need for explicit mathematical modeling. Despite their initial promise, these methods are inherently constrained by their dependence on relatively compact neural networks. Such models exhibit limited representational capacity, which hinders their ability to learn complex, high-dimensional feature representations. Consequently, they are particularly susceptible to performance degradation in the face of diverse and nonstationary environmental conditions. These challenges motivate us to first explore the use of PLMs for radar target detection.
@@ II-B Language Model Powered Signal Processing Task
Recently, LLMs such as ChatGPT and DeepSeek have demonstrated exceptional capabilities in natural language understanding, code generation, math problem solving, and human-like text generation [30]. Some studies also demonstrate their potential in the signal processing (SP) task. For example, Zhou et al. [18] propose a unified time series analysis framework by fine-tuning frozen GPT2 [31] and achieving SOTA performance on various datasets. Liu et al. [24] propose a GPT2-empowered channel prediction framework to improve prediction accuracy. Sheng et al. [32] utilize the GPT2 model to develop an effective and robust beam prediction method. Zheng and Dai [21] propose a multi-task PLM framework to satisfy the requirements of different wireless communication tasks. IOT-LLM and Penetrative AI [33, 34] directly utilize raw signal data as input and chain of thought prompts for reasoning on the Internet of Things (IOT) tasks. However, recent studies [35] have raised concerns about the universal applicability of PLM in signal processing tasks. PLM-based approaches do not always deliver significant performance improvements, implying that their pre-trained parameters sometimes fail to transfer to downstream tasks. In marine radar target detection under low SCR conditions, directly fine-tuning a PLM often leads to severe overfitting issues, limiting its parameter transfer capability. To mitigate this limitation, we introduce a preference-aware loss that selectively trains on informative feature patches.
@@ III Proposed Method
Figure 1: Overview of our RadarPLM framework.
@@
Figure 3: An illustration of the patching operation.
III-A Overview of the RadarPLM Framework
The overall framework of RadarPLM is illustrated in Fig. 1, which consists of four tightly connected stages: (1) sequence feature extraction and patching, (2) reference model training with token-level loss computation, (3) fine-tune LLMs for marine radar target detection, and (4) autoencoder-based binary classification head retraining.
To enable lightweight and generalizable fine-tuning of PLMs for radar target detection, the proposed RadarPLM framework integrates four tightly coupled modules into a unified optimization pipeline. First, the sequence feature extraction and patching module (Section III-B) derives five discriminative sequence features from the multi-domain transformation of radar echo signals. Then a patching module is devised to capture local semantic patterns while reducing computational complexity. Second, a lightweight adaptation module (Section III-C) is introduced to efficiently adapt the PLM backbone to the radar task, retaining its universal knowledge while minimizing parameter updates. Third, a preference-aware loss function (Section III-D), leveraging a lightweight reference model, enables selective optimization and mitigates overfitting. As shown in Fig. 1, the PLM-based model utilizes an input embedding layer to project feature patches into the PLM’s encoder feature space. The embedded representations are processed by the initial encoder layers to extract more discriminative features. A layer normalization operation refines these features, producing the normalized representation $F^{\mathrm{LN}}$, which enhances consistency and stability. Finally, two output linear layers generate token-level detection results. For the reference model structure, a standard Transformer encoder [36] processes the input feature patches to produce token-level detection outputs, as illustrated in Fig. 2. An autoencoder-based binary classification head (Section III-E) is retrained to further improve detection rate.
@@ III-B Sequence Feature Extraction and Patching
III-C Lightweight Fine-tuning Module for the PLM Backbone
Adapting large-scale PLMs to specialized downstream applications like radar target detection presents significant challenges, necessitating a lightweight fine-tuning module. This necessity stems from two primary considerations. First, the immense parameter scale of PLMs renders full fine-tuning computationally expensive and often impractical for resource-constrained environments. Second, and more critically for radar target detection, available datasets are typically limited in size. Fine-tuning an entire model on such small datasets introduces the risks of catastrophic forgetting, where the model’s valuable pre-trained knowledge is overwritten, and overfitting, due to data scarcity.
Drawing inspiration from recent work [18, 37], we propose a lightweight fine-tuning module designed to balance this trade-off between adaptability and efficiency. Our approach is guided by two hypotheses, consistent with [18, 37]: (1) achieving strong performance does not require updating all network parameters; instead, updating a small subset, such as bias terms or specific layers, while keeping the majority frozen is sufficient. (2) The weight updates during adaptation possess a low ”intrinsic rank.” Consequently, instead of modifying the entire weight matrix, we can capture task-specific knowledge within a much smaller, low-rank subspace. Our module implements these insights by jointly updating the layer normalization parameters and incorporating a Low-Rank Adaptation (LoRA) component.
Concretely, we apply LoRA to the query and value projection matrices, $\mathbf{W}_{Q}$ and $\mathbf{W}_{V}$, within each attention block of the PLM backbone, as illustrated in Fig. 5. For a given frozen weight matrix $\mathbf{W}$, LoRA introduces two low-rank trainable matrices, $A$ and $B$, and modifies the transformation as follows:
@@ III-C Lightweight Fine-tuning Module for the PLM Backbone
where $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$, $A\in\mathbb{R}^{d_{1}\times r}$, $B\in\mathbb{R}^{d_{2}\times r}$, and $r\ll\min(d_{1},d_{2})$. The low-rank decomposition effectively projects the task-specific adaptation into a compact subspace, greatly reducing the parameter footprint while retaining original pre-trained knowledge. The matrices are initialized as $A\sim\mathcal{N}(0,\sigma^{2})$ and $B=0$, ensuring stable convergence and preventing interference with pre-trained weights during the early training phase. During fine-tuning, only $A$ and $B$ are updated, while $\mathbf{W}$ remains fixed. Owing to the low-rank constraint: $r\ll\min(d_{1},d_{2})$, the number of parameters for fine-tuning $A$ and $B$, i.e., $rd_{1}+rd_{2}$, is significantly less than that of the full weight matrix, $d_{1}d_{2}$. The reduced parameter size thus makes the fine-tuning much efficient.
This design offers two key advantages. First, this module mitigates the computationally expensive limitation of full fine-tuning by updating only a very small subset of the network parameters and compact low-rank matrices, thereby reducing trainable parameters by several orders of magnitude. Second, this module alleviates catastrophic forgetting of pre-trained knowledge during adaptation, as the newly acquired information is encoded within a subspace rather than the original representation. Consequently, this module enables the finetuned model to retain its pre-trained general-purpose knowledge while efficiently adapting to downstream radar tasks.
@@ III-D1 Methodology
To address the prevalent issue of overfitting in PLM-based model fine-tuning, we introduce a preference-aware loss function to enable selective training, a novel strategy designed to guide the model towards learning generalizable feature patterns. During fine-tuning, feature patches dominated by clutter-induced noise or anomalies can mislead the model and hinder effective learning. Conventional optimization methods, which treat all feature patterns indiscriminately with a standard cross-entropy loss, are prone to memorizing these non-generalizable, instance-specific characteristics, leading to severe overfitting.
In contrast, our proposed selective training strategy follows a ”preference-aware” philosophy. The core idea is to evaluate the learning value of each feature patch during training and use this evaluation to reweight its contribution to the final loss. This approach enables the model to focus on feature patterns with high learning values while minimizing the impact of noisy or anomalous patterns. By preventing the model from overfitting to these non-generalizable patterns, this strategy enhances model robustness and mitigates overfitting.
Specifically, for each feature token $F_{b,k}^{\mathrm{fi}}$ within a training batch, we compute two distinct loss values:
  • Reference Loss $\mathcal{L}_{\theta_{{r}},b,k}(F^{\mathrm{fi}}_{b,k})$: This loss is computed using a lightweight reference model $\theta_{r}$, which has been trained on the validation set. It serves as a stable proxy for evaluating the feature patch’s inherent difficulty or noise level:
@@
    where $\hat{y}_{b,k}^{i,\theta_{r}}$ is the prediction from lightweight reference model $\theta_{r}$.
  • Target Loss $\mathcal{L}_{\theta_{{t}},b,k}(F^{\mathrm{fi}}_{b,k})$: his loss is calculated using the predictions from the current PLM-based model $\theta_{t}$ being fine-tuned, which reflects how well the model currently fits the feature patch:
@@
    where $\hat{y}_{b,k}^{i,\theta_{t}}$ represents the prediction from the current PLM-based model $\theta_{t}$.
The key to our method is computing an evaluated learning value $s_{b,k}$ based on the discrepancy between the target and reference losses:
@@
where $\alpha$ is a scaling coefficient, and the $\mathrm{ReLU}(\cdot)$ function, defined as $\mathrm{ReLU}(x)=\max(0,x)$, ensures that only tokens with positive excess loss are fine-tuned.
This loss difference naturally leads to a selective training scheme with desirable properties:
  • High Importance (Unlearned): If $\mathcal{L}_{\theta_{{t}}}$ is much larger than $\alpha\mathcal{L}_{\theta_{{t}}}$, the feature token is identified as containing valuable, yet unlearned patterns.
  • Low Importance (Noise/Anomaly): If $\alpha\mathcal{L}_{\theta_{{r}}}$ is large (indicating with high inherent noise), it effectively dampens $s_{b,k}$, even if $\mathcal{L}_{\theta_{{t}}}$ is large, preventing the model from overfitting to noisy or anomalous data.
  • Low Importance (Well Learned): If $\mathcal{L}_{\theta_{{t}}}$ is already small, the feature token is considered well learned, and $s_{b,k}$ will also be small, preventing redundant updates to easy patterns.